# Domain Generalization: A Tale of Two ERMs

### v1

@@ 5.1 Decision-theoretic Gain Under Posterior Drift
Theorem 1 holds regardless of the distribution $P_{XYMD}$. By considering assumptions on this distribution, stronger conclusions may be drawn. We now examine a class of distributions where the gap $R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}$ has a more concrete lower bound. This class of distributions is motivated by applications—particularly in natural language processing—where the posterior $P_{Y\mid X,D}$ (correspondingly, $P_{Y\mid X,M}$) differs across domains due to inherent ambiguity or subjectivity. A canonical example is the sentiment or toxicity annotation task, where annotators often disagree on the same text. For instance, in the age-related sentiment analysis dataset of Díaz et al. (2018), the sentence “Old people’s appearance contains so much lived life.” received conflicting labels: 2/5 annotators seeing it as ‘very positive’, 2/5 as ‘somewhat positive’, and 1/5 as ‘very negative’. This reflects how labeling tendency varies with annotator identity. We capture this phenomenon by introducing a formal posterior drift class for domain generalization.
Definition 2 (Posterior Drift Class for DG)
@@
$P_{XMM^{\prime}}\Big(f^{*}_{\mathrm{DG}}\big(X,M\big)\neq f^{*}_{\mathrm{DG}}\big(X,M^{\prime}\big)\Big)\geq\epsilon\Big\},$
@@
Proposition 2 (Gain Under Posterior Drift)
$\inf_{P_{XYMD}\in\Pi(\gamma,\epsilon)}\Big[R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}\Big]\geq\frac{\gamma\cdot\epsilon}{2}.$
The proof is given in Section˜C.3. This result establishes that, under the posterior drift class $\Pi(\gamma,\epsilon)$, the benefit of leveraging domain-specific information is guaranteed to be at least $\gamma\epsilon/2$. In other words, the risk reduction from domain-informed prediction is not only possible but provably guaranteed in this regime. This proposition can be viewed as a concrete version of the more general lower bound in Theorem˜1.
Remark 4
In contrast to pessimistic results in domain adaptation—where no method consistently outperforms vanilla ERM under posterior drift (Zhu et al., 2024; Wang et al., 2025)—our work presents an optimistic view in domain generalization: by conditioning on domain metadata $M$, we can provably do better than pooling-based prediction.
@@
Theorem 2 ($R^{*}_{\mathrm{DG}}$ vs. $R^{*}_{\mathrm{full}}$)
@@
$\mathbb{E}_{X,D,M}\left[\widetilde{\gamma}(X,D)\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
$\quad\leq R^{*}_{\mathrm{DG}}-R^{*}_{\mathrm{full}}$
$\quad\leq\mathbb{E}_{X,D,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DG}}(X,M)}\right].$
The result parallels Theorem˜1: if domain information is low quality—i.e., $f^{*}_{\mathrm{full}}$ often disagrees with $f^{*}_{\mathrm{DG}}$ with high-confidence—then $R^{*}_{\mathrm{DG}}$ can be much larger than $R^{*}_{\mathrm{full}}$. The necessary and sufficient conditions for this gap to vanish are analogous to those in Theorem˜1.
5.3 Gain Under Covariate Shift?
Covariate shift refers to a setting where, as the domain information $M$ varies, $P_{X|M}$ changes, but $P_{Y|X,M}$ does not (Quiñonero-Candela et al., 2008). More generally, we may extend the meaning of covariate shift to be any DG problem where the Bayes classifier $f^{*}_{\mathrm{DG}}$ does not depend on $M$. In such a scenario, access to domain-specific information cannot reduce the Bayes risk.
Corollary 1
Under covariate shift, $R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}$.
Nevertheless, empirical studies have reported performance gains from incorporating domain information even under covariate shift (Dubey et al., 2021; Bui et al., 2021). Our next result shows that such gains do not reflect any decision-theoretic benefit, but instead arise from the restrictions of the chosen function class.
Assumption 1 (DI & Pooling Function Classes)
@@ 5.3 Gain Under Covariate Shift?
    For every $g\in\mathcal{G}_{k}$, there exists $f\in\mathcal{F}_{k}$ and some fixed $m_{0}\in\mathcal{M}$ such that
    $g(x)=f(x,m_{0}).$
@@
The first condition is satisfied by common architectures (e.g., feedforward ReLU networks) and $\mathcal{G}$ can be induced from $\mathcal{F}$ by setting the weights from the $m$ inputs to zero. What’s more, $\mathcal{G}$ satisfies the universal approximation property as the network size grows.
Proposition 3
@@
$R^{*}_{\mathrm{pool},\mathcal{G}_{k}}:=\inf_{g\in\mathcal{G}_{k}}R(g)\ \text{ and }\ R^{*}_{\mathrm{DG},\mathcal{F}_{k}}:=\inf_{f\in\mathcal{F}_{k}}R(f).$
Under covariate shift,
$R^{*}_{\mathrm{pool},\mathcal{G}_{k}}\geq R^{*}_{\mathrm{DG},\mathcal{F}_{k}}\quad\forall k,$ (4)
@@
$\lim_{k\to\infty}R_{\mathrm{pool},\mathcal{G}_{k}}=\lim_{k\to\infty}R_{\mathrm{DG},\mathcal{F}_{k}}.$ (5)
This shows that with a restricted function class $\mathcal{G}_{k}$, leveraging domain-information may lead to a benefit under covariate shift. Furthermore, strict inequality in Equation˜4 is possible, as illustrated with a simple example in Appendix˜B. However, these gains vanish asymptotically as the function classes become more expressive. This phenomenon is also observed empirically in Section˜6.
fnum@section6 Experiments
We evaluate the effectiveness of domain-informed ERM (DI-ERM) in three experimental settings. Our primary focus is on the comparison between DI-ERM and pooling ERM, which highlights the benefit of incorporating domain metadata. Additional results—including linear probing, benchmarks against alternative and state-of-the-art methods, and complete experimental details—are provided in Appendix˜D.
@@ 6.1 Annotator Disagreement
In many NLP tasks, annotators exhibit subjective preferences, leading to disagreement on the label $y$ for the same input xx—a form of posterior drift discussed in Section˜5.1. To study this phenomenon, we use the dataset of Díaz et al. (2018), which re-annotates a subset of the Sentiment140 dataset (Go et al., 2009) for training and provides a test set drawn from age-related blog posts. The training set comprises 59,235 sentences labeled by 1,481 annotators; the test set includes 1,419 sentences labeled by 878 annotators. Each sentence is annotated by 4–5 individuals, and the labels exhibit high disagreement (about 40 $\%$). In this setting, the input $x$ is a sentence, the label $y\in\{1,2,3,4,5\}$ denotes sentiment on a five-point scale, the domain $d$ corresponds to the annotator, and the domain information $m$ consists of demographic metadata (e.g., age, upbringing region).
To encode domain information $M$, we concatenate it with the sentence $x$ in a text-prompt format, as illustrated in Figure˜3. We finetune the BERT base model (Devlin et al., 2019). Table˜2 reports the results. Since prior DG methods have not been evaluated on this dataset, we implemented and benchmarked them under the same experimental setup. DI-ERM substantially outperforms pooling ERM, demonstrating that leveraging annotator metadata can dramatically improve predictive accuracy.
Table 2: Test accuracy on the sentiment disagreement dataset. Previous DG methods had not been benchmarked on this dataset; we re-implemented them for comparison under our setup. Incorporating annotator profiles ($M$) through DI-ERM yields a dramatic improvement over pooling ERM, reflecting the importance of modeling annotator-specific posterior drift. In particular, DI-ERM nearly doubles accuracy compared to pooling ERM and surpasses the previous state-of-the-art by Deng et al. (2023).
@@
AnnEmb (SOTA) BERT 64.6 $\pm$ 0.8
DI-ERM (ours) BERT 90.5 $\pm$ 0.2
@@ 6.2 Reviewer-specific Analysis
We next examine the WILDS-Amazon Reviews dataset (Koh et al., 2021), which captures distributional shifts across reviewers. Here, the input $x$ is a product review, $y\in\{1,\dots,5\}$ is the star rating, $d$ denotes the reviewer identity, and $m$ consists of all (unlabeled) reviews written by that reviewer. Once again, because of differences among reviewers, posterior drift is exhibited.
@@
To incorporate reviewer context $M$, we randomly sample 20 additional reviews written by the same reviewer and concatenate them with the current review in a prompt format, shown in Figure˜4. We finetune the nomic-embed-text-v1.5 model (Nussbaum et al., 2025), which supports a longer context window than BERT so that domain information $M$ can fit in. As summarized in Table˜3, DI-ERM outperforms pooling ERM. Beyond higher average accuracy, DI-ERM also boosts the 10th-percentile accuracy across reviewers—a key robustness metric used on the official leaderboard.
[…]
@@
Pooling ERM nomic-embed 72.2 $\pm$ 1.1 55.1 $\pm$ 0.8
DI-ERM (ours) nomic-embed 73.4 $\pm$ 0.1 56.4 $\pm$ 0.8
@@ 6.3 Image Classification Across Styles
We next evaluate our method on the PACS dataset (Li et al., 2017), which contains images from four distinct visual styles (aka, domain $d$): Photo (P), Art Painting (A), Cartoon (C), and Sketch (S). Each image $x$ belongs to one of seven categories: dog, elephant, giraffe, guitar, horse, house, or person. Domain information $m$ is encoded by a short text description $m$, such as “a photo” or “a pencil sketch” (see Figure˜5).
This vision task satisfies covariate shift, since a single classifier should accurately classify all images across domains. Thus, in line with Section˜5.3, we expect any gains to be due to using a restricted function class. To implement DI-ERM, we use pretrained image foundation models (e.g., CLIP (Radford et al., 2021), DINOv2 (Oquab et al., 2023), DINOv3 (Siméoni et al., 2025)) to extract visual features from $x$, and encode the domain description $m$ using a pretrained language model (DistilBERT) following the prompt in Figure˜5. The resulting image and text embeddings are concatenated into a joint representation for classification.
We follow the standard PACS evaluation protocol: training on three domains and testing on the held-out fourth domain, repeated across all domain splits. All encoders are frozen, and linear classifiers are trained on top of the fixed representations.
As shown in Section˜6.3, DI-ERM improves over pooling ERM in most settings. The gains are most pronounced for mid-sized models, while the benefit diminishes for larger foundation models. This pattern aligns with the discussion in Section˜5.3: under covariate shift, the benefit of DI-ERM decreases as model size grows.
Table 4: Domain generalization results on PACS using models from the CLIP, DINOv2 and DINOv3 families. DI-ERM achieves improved accuracy over pooling ERM in most configurations, moreso for mid-sized models than for large models, a phenomenon consistent with our theory in Section˜5.3. When using large models, both ERM and DI-ERM approach SOTA performance by Cho et al. (2023).
@@
Pooling ERM (linear) CLIP: ViT-small 95.0 $\pm$ 0.0
@@
gray!80blackPooling ERM (linear) CLIP: ViT-large 98.6 $\pm$ 0.0
@@
Pooling ERM (linear) DINOv2: ViT-small 87.6 $\pm$ 0.0
@@
gray!80blackPooling ERM (linear) DINOv2: ViT-large 96.1 $\pm$ 0.0
@@
Pooling ERM (linear) DINOv3: ViT-small 91.7 $\pm$ 0.0
@@
gray!80blackPooling ERM (linear) DINOv3: ViT-large 95.4 $\pm$ 0.0
@@
fnum@section7 Conclusions
This work presents a rigorous theory of domain generalization, precisely characterizing when and why leveraging domain information at test time is beneficial. Empirically, we demonstrate that domain-informed ERM (DI-ERM) outperforms pooled ERM across three representative scenarios in language and vision tasks. A key requirement of DI-ERM is access to domain metadata; when such metadata is unavailable, unlabeled data from the target domain offers a natural alternative.
Future work might explore alternative ways of encoding domain information, and a broader range of DG benchmarks. Additionally, it would interesting to explore the phenomena discussed in the context of other performance measures for DG, such as the worst-case test error examined in (Dwork et al., 2025).

### v2

@@ 5.1 Decision-Theoretic Gain Under Posterior Drift
Theorem 1 holds regardless of the distribution $P_{XYMD}$. By imposing structural assumptions, stronger conclusions can be drawn. We now examine a class of distributions where the gap $R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DI}}$ has a more concrete lower bound. This class of distributions is motivated by applications, particularly in NLP, where the posterior $P_{Y\mid X,D}$ (correspondingly, $P_{Y\mid X,M}$) differs across domains due to inherent ambiguity or subjectivity. A canonical example is the sentiment or toxicity annotation task, where annotators often disagree on the same text. For instance, in the age-related sentiment analysis dataset of Díaz et al. (2018), the sentence “Old people’s appearance contains so much lived life.” received conflicting labels: 2/5 annotators seeing it as ‘very positive’, 2/5 as ‘somewhat positive’, and 1/5 as ‘very negative’. This reflects how labeling tendency varies with annotator identity (aka, domain). We capture this phenomenon by introducing a formal posterior drift class for DG.
Definition 2 (Posterior Drift for DG).
For $\gamma,\epsilon>0$, denote
@@
$P_{XMM^{\prime}}\Big(f^{*}_{\mathrm{DI}}\big(X,M\big)\neq f^{*}_{\mathrm{DI}}\big(X,M^{\prime}\big)\Big)\geq\epsilon\Big\},$
@@
Theorem 2 (Gain Under Posterior Drift).
$\inf_{P_{XYMD}\in\Pi(\gamma,\epsilon)}\Big[R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DI}}\Big]\geq\frac{\gamma\cdot\epsilon}{2}.$
The proof is given in Section C.3. This result establishes that, under the posterior drift class $\Pi(\gamma,\epsilon)$, the benefit of leveraging domain-specific information is guaranteed to be at least $\gamma\epsilon/2$. This result can be viewed as a concrete version of the more general lower bound in Theorem 1.
Remark 4.
In contrast to pessimistic results in domain adaptation, where no method consistently outperforms vanilla ERM under posterior drift (Zhu et al., 2024; Wang et al., 2025), our work presents an optimistic view in domain generalization: by conditioning on domain metadata $M$, we can provably do better than pooling-based prediction.
@@
Theorem 3 ($R^{*}_{\mathrm{DI}}$ vs. $R^{*}_{\mathrm{full}}$).
@@
$\mathbb{E}_{X,D,M}\left[\widetilde{\gamma}(X,D)\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]$
$\quad\leq R^{*}_{\mathrm{DI}}-R^{*}_{\mathrm{full}}$
$\quad\leq\mathbb{E}_{X,D,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DI}}(X,M)}\right].$
The result parallels Theorem 1: if domain information is low quality, i.e., $f^{*}_{\mathrm{full}}$ often disagrees with $f^{*}_{\mathrm{DI}}$ with high-confidence, then $R^{*}_{\mathrm{DI}}$ can be much larger than $R^{*}_{\mathrm{full}}$. The necessary and sufficient conditions for this gap to vanish are analogous to those in Theorem 1.
5.3 Gain Under UBC?
The preceding subsection establishes that domain information can strictly improve performance under posterior drift. We now contrast this with the universal Bayes classifier (UBC) regime, in which a single function of the input $X$ realizes the Bayes classifier for all domains.
Definition 3 (Universal Bayes Classifier).
A DG problem satisfies universal Bayes classifier assumption if
$f^{*}_{\mathrm{pool}}(x)=f^{*}_{\mathrm{DG}}(x,m)=f^{*}_{\mathrm{full}}(x,d)\quad\forall x,m,d.$
This is inspired by applications in vision where a single classifier performs well across all domains. A concrete example of this is when the support of $P_{X|D}$ is disjoint across domains. The next result follows from Theorem 1.
Corollary 1.
Under UBC, $R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DI}}$.
In such a scenario, access to domain-specific information cannot reduce the Bayes risk. Nevertheless, studies have reported empirical performance gains from incorporating domain information even under UBC (Dubey et al., 2021; Bui et al., 2021). Our next result shows that such gains arise from the choice of function classes.
Assumption 1 (DI & Pooling Function Classes).
@@ A DG problem satisfies universal Bayes classifier assumption if
    For every $g\in\mathcal{G}_{k}$, there exists $f\in\mathcal{F}_{k}$ such that
    $g(x)=f(x,m)\quad\forall m\in\mathcal{M}.$
@@
The first condition requires that the pooling class $\mathcal{G}_{k}$ can be represented by the domain-informed class $\mathcal{F}_{k}$. This condition is satisfied by common architectures such as feedforward ReLU networks, where $\mathcal{G}_{k}$ can be induced from $\mathcal{F}_{k}$ by setting the weights associated with the metadata inputs to zero. The second condition assumes that the pooling classes $\{\mathcal{G}_{k}\}$ are sufficiently expressive to approximate $f^{*}_{\mathrm{pool}}$.
Proposition 2.
@@
$R^{*}_{\mathrm{pool},\mathcal{G}_{k}}:=\inf_{g\in\mathcal{G}_{k}}R(g)\ \text{ and }\ R^{*}_{\mathrm{DI},\mathcal{F}_{k}}:=\inf_{f\in\mathcal{F}_{k}}R(f).$
Under UBC,
$R^{*}_{\mathrm{pool},\mathcal{G}_{k}}\geq R^{*}_{\mathrm{DI},\mathcal{F}_{k}}\quad\forall k,$ (4)
@@
$\lim_{k\to\infty}R^{*}_{\mathrm{pool},\mathcal{G}_{k}}=\lim_{k\to\infty}R^{*}_{\mathrm{DI},\mathcal{F}_{k}}.$ (5)
This shows that with a restricted function class $\mathcal{G}_{k}$, leveraging domain-information may lead to a benefit under UBC. Indeed, strict inequality in Equation 4 is possible, as illustrated with a simple example in Appendix B. However, these gains vanish asymptotically as the function classes become more expressive. This phenomenon is also observed empirically in Section 6.
6 Experiments
We evaluate the effectiveness of domain-informed ERM (DI-ERM) in three experimental settings. The goal is not to achieve state-of-the-art performance, but to empirically validate the theoretical insights from previous sections regarding the benefits of leveraging domain metadata under posterior drift (e.g., Theorems 1 and 2).
Methods
Our primary comparison is between DI-ERM and pooling-ERM. This comparison directly tests our central theoretical claim: incorporating domain metadata at inference time can strictly improve performance under posterior drift, while pooling-ERM is suboptimal.
We also include several widely used DG baselines: IRM (Arjovsky et al., 2019), GroupDRO (Sagawa et al., 2020), and CORAL (Sun and Saenko, 2016). These methods exploit domain indices during training but make predictions using only the input $X$ at inference time, and therefore fall within the domain-agnostic paradigm.
Closest to our setting, D3G (Yao et al., 2024) also leverages domain metadata at both training and inference time. Unlike DI-ERM, D3G trains domain-specific prediction heads—one for each training domain—and uses metadata to reweight their predictions during inference. This design can become less robust when the number of training domains is large relative to the amount of labeled data per domain. In contrast, DI-ERM encodes domain metadata directly as additional input, yielding a simpler and empirically more effective approach in such regimes.
Additional results, including linear probing, benchmarks against alternative and state-of-the-art methods, and complete experimental details, are provided in Appendix D.
@@ 6.1 Annotator Disagreement
In many NLP tasks, annotators exhibit subjective preferences, leading to disagreement on the label $y$ for the same input xx—a form of posterior drift discussed in Section 5.1. To study this phenomenon, we use the dataset of Díaz et al. (2018), which re-annotates a subset of the Sentiment140 dataset (Go et al., 2009) for training and provides a test set drawn from age-related blog posts. The training set comprises 59,235 sentences labeled by 1,481 annotators; the test set includes 1,419 sentences labeled by 878 annotators. Each sentence is annotated by 4–5 individuals, and the labels exhibit high disagreement (about 40 $\%$), a sign of large posterior drift. In this setting, the input $x$ is a sentence, the label $y\in\{1,2,3,4,5\}$ denotes sentiment on a five-point scale, the domain $d$ corresponds to an annotator, and the domain information $m$ consists of demographic metadata (e.g., age, upbringing region).
To encode domain information $M$, we concatenate it with the sentence $x$ in a text-prompt format (Figure 3 in Appendix D). We finetune the BERT base model (Devlin et al., 2019). Table 2 reports the results. Since prior DG methods have not been evaluated on this dataset, we implemented and benchmarked them under the same experimental setup. DI-ERM substantially outperforms pooling-ERM, demonstrating that leveraging annotator metadata can dramatically improve predictive accuracy under posterior drift.
Table 2: Test accuracy on the sentiment disagreement dataset. Previous DG methods had not been benchmarked on this dataset; we re-implemented them for comparison under our setup. Incorporating annotator profiles ($M$) through DI-ERM yields a dramatic improvement over pooling-ERM, reflecting the importance of modeling annotator-specific posterior drift.
@@
Pooling-ERM BERT 49.1 $\pm$ 0.4
D3G BERT 54.2 $\pm$ 0.4
DI-ERM BERT 90.5 $\pm$ 0.2
@@ 6.2 Reviewer-specific Analysis
We next examine the WILDS-Amazon Reviews dataset (Koh et al., 2021), which captures distributional shifts across reviewers. Here, the input $x$ is a product review, $y\in\{1,\dots,5\}$ is the star rating, $d$ denotes the reviewer identity, and $m$ consists of all (unlabeled) reviews written by that reviewer. Once again, because of differences among the rating patterns of reviewers, posterior drift is exhibited.
@@
To incorporate reviewer context $M$, we randomly sample 20 additional reviews written by the same reviewer and concatenate them with the current review in a prompt format (Figure 4 in Appendix D). We finetune the nomic-embed-text-v1.5 model (Nussbaum et al., 2025), which supports a longer context window than BERT so that domain information $M$ can fit in. As summarized in Table 3, DI-ERM outperforms pooling-ERM. Beyond higher average accuracy, DI-ERM also boosts the 10th-percentile accuracy across reviewers—a key robustness metric used on the official leaderboard.
Table 3: Sentiment classification performance on Amazon-WILDS with reviewer-specific context. The first block reports results quoted directly from the official WILDS leaderboard (https://wilds.stanford.edu/), which uses DistilBERT. The second block reports our own experiments using nomic-embed-text-v1.5 model. DI-ERM improves over pooling-ERM, both in average accuracy and in 10th-percentile reviewer accuracy—the official leaderboard metric.
@@
Pooling-ERM nomic-embed 72.2 $\pm$ 1.1 55.1 $\pm$ 0.8
D3G nomic-embed 71.8 $\pm$ 0.1 54.7 $\pm$ 0.0
DI-ERM nomic-embed 73.4 $\pm$ 0.1 56.4 $\pm$ 0.8
@@ 6.3 Image Classification Across Styles
Lastly, we evaluate on a setup where there is little to no posterior drift. This is an empirical verification of the theory in Section 5.3. We choose PACS dataset (Li et al., 2017), which contains images from four distinct visual styles (aka, domain $d$): Photo (P), Art Painting (A), Cartoon (C), and Sketch (S). Each image $x$ belongs to one of the seven categories: dog, elephant, giraffe, guitar, horse, house, or person. Domain information $m$ is encoded by a short text description $m$, such as “a photo” or “a pencil sketch” (Figure 5 in Appendix D).
This vision task satisfies UBC, since a single classifier based on image $x$ should accurately classify all images across domains. Thus, in line with Section 5.3, we expect any gains from domain metadata to be due to using a restricted function class. To implement DI-ERM, we use pretrained image foundation models (e.g., CLIP (Radford et al., 2021), DINOv2 (Oquab et al., 2023), DINOv3 (Siméoni et al., 2025)) to extract visual features from $x$, and encode the domain description $m$ using a pretrained language model (DistilBERT) following the prompt in Figure 5. The resulting image and text embeddings are concatenated into a joint representation for classification.
We follow the standard PACS evaluation protocol: training on three domains and testing on the held-out fourth domain, repeated across all domain splits. All encoders are frozen, and linear classifiers are trained on top of that.
As shown in Table 4, the gains of DI-ERM over pooling-ERM are most pronounced for ViT-small models, while the benefit diminishes for ViT-large foundation models. This pattern aligns with the discussion in Section 5.3: under UBC, the benefit of DI-ERM decreases as model size grows.
Table 4: Domain generalization results on PACS using models from the CLIP, DINOv2 and DINOv3 families. DI-ERM achieves improved accuracy over pooling-ERM, but the gain diminishes for large models, a phenomenon consistent with our theory in Section 5.3. When using large models, both ERM and DI-ERM approach SOTA performance by Cho et al. (2023).
@@
Pooling-ERM (linear) CLIP: ViT-small 95.0 $\pm$ 0.0
@@
Pooling-ERM (linear) CLIP: ViT-large 98.6 $\pm$ 0.0
@@
Pooling-ERM (linear) DINOv2: ViT-small 87.6 $\pm$ 0.0
@@
Pooling-ERM (linear) DINOv2: ViT-large 96.1 $\pm$ 0.0
@@
Pooling-ERM (linear) DINOv3: ViT-small 91.7 $\pm$ 0.0
@@
Pooling-ERM (linear) DINOv3: ViT-large 95.4 $\pm$ 0.0
@@
7 Conclusions
This work presents a rigorous theory of domain generalization, and identifies posterior drift as the key regime in which domain information at test time can be strictly beneficial. Empirically, we demonstrate that domain-informed ERM (DI-ERM) outperforms pooling-ERM across several representative scenarios under posterior drift.
A central practical requirement of DI-ERM is access to domain metadata; when such metadata is unavailable, unlabeled data from the target domain offers a natural alternative. Future work might explore alternative ways of encoding domain information, and a broader range of DG benchmarks. Additionally, it would be interesting to explore other performance measures for DG, such as the worst-case test error in Dwork et al. (2025).